the sum of the square of Two numbers Is 225 and the square of their sum is 441

What is the reciprocal of their product


1
Expert's answer
2020-08-10T14:07:49-0400

Let the two numbers be x and y

( x + y )2 = 441

x2 + y2 = 225

Therefore;

( x + y )2 - ( x2 + y2 ) = 441 - 225

( x + y)2 = ( x + y )(x + y)

= x( x+y) + y(x+y)

= x2 +xy +xy +y2

= x2 + 2xy + y2

Therefore;

(x2 + 2xy + y2) - ( x2 + y2 ) = 441 - 225

2xy/2 = 216/2

xy = 108

The reciprocal of xy is 1/108